OpenStudy (anonymous):

A glacier moves about 12 inches every 36 hours. About how far does the glacier move in one week?

OpenStudy (anonymous):

Hey, a good way to do this problem is by finding out how many inches does it move in 24 hours.

OpenStudy (anonymous):

168/36=4.6X12=56. Correct me if im wrong

OpenStudy (anonymous):

Or you can set it up as so, \[\frac{ 12 inches }{ 36 hours } = \frac{ x }{ }\]

OpenStudy (anonymous):

x/ 168 hours

OpenStudy (anonymous):

I did that but I didn't know what to do from there

OpenStudy (anonymous):

\[\frac{ 12inches }{ 36h } = \frac{ x }{ 168h }\]

OpenStudy (anonymous):

Oh you have to cross multiply

OpenStudy (anonymous):